OpenStudy (anonymous):

If the triangles are similar, which ratio shows the relationship of the area of the larger triangle to the area of the smaller triangle?

OpenStudy (anonymous):

I think it's 6:1

OpenStudy (callisto):

Nope... for similar figures, ratio of area = (ratio of sides)^2 area of the larger triangle : the area of the smaller triangle = ( side of the larger triangle : the side of the smaller triangle)^2 = (6/1)^2 = 36 /1 = 36:1 Got it?

OpenStudy (anonymous):

Oh yes thanks :D

OpenStudy (callisto):

welcome :)